OpenStudy (anonymous):

Trying to figure out if I did this correctly. If someone could look at my work it would be appreciated! :) Indefinite integral of dx/x^2 (sqrt(1+x^2) (I'll put my work in the comments section)

OpenStudy (anonymous):

\[\int\frac{dx}{x^2\sqrt{1+x^2}}~~?\]

OpenStudy (anonymous):

If \[a^2+u^2\] then u=a tan theta and \[1 + \tan \theta^2=\sec \theta^2\] a=1 and u=x x=tan theta and dx=sec^2theta theta=arctan x so \[\int\limits (\sec \theta^2 d \theta)/\tan \theta^2 \sqrt (1+\tan \theta^2) \] and then I ended up with \[\int\limits 1+\tan \theta^2/ \tan \theta^2 (1+\tan \theta)\] I cancelled out some things (not sure if I was supposed to) and ended up with \[\int\limits 1/\tan \theta\] and that is equal to \[1/\sec \theta^2\]

OpenStudy (anonymous):

yes, that's the correct equation, SithsAndGiggles

OpenStudy (anonymous):

Your substitution is right, but I get lost with what you do after that. \[\begin{align*}\int\frac{\sec^2\theta}{\tan^2\theta\sqrt{1+\tan^2\theta}}~d\theta&=\int\frac{1+\tan^2\theta}{\tan^2\theta\sqrt{1+\tan^2\theta}}~d\theta\\ &=\int\frac{\sqrt{1+\tan^2\theta}}{\tan^2\theta}~d\theta\\ &=\int\frac{\sqrt{\sec^2\theta}}{\tan^2\theta}~d\theta\\ &=\int\frac{\sec\theta}{\tan^2\theta}~d\theta\\ &=\int\frac{\cos\theta}{\sin^2\theta}~d\theta \end{align*}\]

OpenStudy (experimentx):

let x=1/t

OpenStudy (anonymous):

so i just realized my integration is NOT correct...but is the work up to that point even close?

OpenStudy (experimentx):

and get \[ I = -\int \frac{t}{\sqrt{t^2 + 1}}dt\]

OpenStudy (anonymous):

^^The alternative method works so much more smoothly :)

OpenStudy (anonymous):

Thank you both!!!

OpenStudy (anonymous):

np